OpenStudy (anonymous):

Solve the following formula for r: C=2πr Choose ALL equivalent variations of your answer. The first answer I got is r=c/2π. What are the other variations?

OpenStudy (anonymous):

@Crixpack please help. Same question as before, except its C=2πr, solve for r. My first answer is r=c/2π.

OpenStudy (anonymous):

That is correct, can you show me your options like you did last time

OpenStudy (anonymous):

@Crixpack I checked off what I think is correct.

OpenStudy (anonymous):

2 of those are incorrect. you have \[r = \frac{ C }{ 2\pi }\] and this is the same as \[r = (\frac{ 1 }{ 2 })(\frac{ C }{ \pi }) = (\frac{ 1 }{ \pi })(\frac{ C }{ 2 })\] etc. What I mean is, look carefully and don't mix up which terms are in the numerator and which are in the denominator.

OpenStudy (anonymous):

@Crixpack, Oh ok, I think I see what I messed up on. is it correct now?

OpenStudy (anonymous):

Yes indeed :)

OpenStudy (anonymous):

Thank you!